OpenStudy (anonymous):

Can someone please explain how to solve 45-45-90 angles, and the theorem used to solve it.

OpenStudy (anonymous):

45-45-90 triangles... in other words, isosceles right triangles :)

OpenStudy (anonymous):



OpenStudy (anonymous):

So, you're usually given one of the sides and asked to determine the rest, right? ^_^

OpenStudy (anonymous):

Yes :) I know the hypotenuse equals 2 squared times the leg, right?

OpenStudy (anonymous):

You mean square root of 2, right? :/

OpenStudy (anonymous):



OpenStudy (anonymous):

Um yes :s I'm confused with all of it really O_o I've been trying to understand it for the past hour

OpenStudy (anonymous):

Well, you have a problem we can work on?

OpenStudy (anonymous):

Yeah, here's one :)

OpenStudy (anonymous):

f=√ ̅2 times 44 would be the first step? O_o

jhonyy9 (jhonyy9):

so you need calculi the side p ?

OpenStudy (anonymous):

yes ._. sorry I meant P* earlier,

jhonyy9 (jhonyy9):

than sin 45 = p/44

jhonyy9 (jhonyy9):

us this

jhonyy9 (jhonyy9):

hope you know the value of sin 45 ,yes ?

jhonyy9 (jhonyy9):

come on please cooperat anything

OpenStudy (anonymous):

Um 0.7071?

jhonyy9 (jhonyy9):

write every calculi please here step by step

jhonyy9 (jhonyy9):

but in the calculi for sin and cos 45 we us sqrt2 / 2

OpenStudy (anonymous):

yeah sorry I'm not following you O.o

jhonyy9 (jhonyy9):

so than sin 45 = p/44 sqrt2 /2 = p/44 p= 44sqrt2/2 = 22sqrt2 p=22sqrt2

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

yw good luck bye

OpenStudy (anonymous):

Well I'm sure I'll understand it sometime. :) Thanks for your help though

jhonyy9 (jhonyy9):

yw was my pleasure